Does number of goof friends depend on happiness? 

(GSS Data)


Very happy                Pretty happy                         Not too happy

Mean              10.4                                        7.4                                          8.3                  

 

Std. dev.        17.8                                        13.6                                        15.6

 

N                    27.6                                        468                                         87

 


 

                                   Sum of                       Mean

 

Source                       squares         df        square           F         Sig

 

Between-groups      1627               2         813                 3.47    0.032

 

Within-groups          193901           828     234

 

Total                          195528           830

1
Expert's answer
2021-12-07T11:24:57-0500

The hypotheses tested are,

"H_0:" Number of good friends depend on happiness

"Against"

"H_1:" Number of good friends does not depend on happiness


We are going to use the given ANOVA table results to make a decision.

From the ANOVA table, the test statistic is "F=3.47" and the p-value "p(F\gt 3.47 )=0.032".

The p-value is compared with the desired significance level of "\alpha=0.05" and the null hypothesis is rejected if "p -value\lt \alpha."

Since "p-value=0.032\lt \alpha=0.05," we reject the null hypothesis and conclude that there is insufficient evidence to show that the number of friends depend on happiness at 5% level of significance.